# High-order Computation of Floquet Multipliers and Subspaces using Multistep Methods

### v1

@@ 2.4 Subspace Perturbation Theory
  3. 3.
    The separation of $A$ and $B$ with respect to the matrix $\mathrm{F}$-norm:
    $\operatorname{sep}_{\mathrm{F}}(A,B)=\inf_{\left\lVert P\right\rVert=1}\left\lVert PA-BP\right\rVert_{\mathrm{F}}.$
Directly computing the separation is hard, but it can be bounded by the spectral gap and departure from normality. In theorem˜1, the upper bound is given in [17], and the lower bound is given by [18].
Theorem 1 (Bounds of $\operatorname{sep}(A,B)$ [18, 17]).
Let $A\in\mathbf{C}^{n\times n}$ and $B\in\mathbf{C}^{m\times m}$ with $\lambda(A)\cap\lambda(B)=\varnothing$. Then we have
$\operatorname{d}(A,B)\geqslant\operatorname{sep}_{\mathrm{F}}(A,B)\geqslant\frac{\operatorname{d}(A,B)}{1+\sum_{j=1}^{dn-2}\left[\frac{\Delta_{\mathrm{F}}(A)+\Delta_{\mathrm{F}}(B)}{\operatorname{d}(A,B)}\right]^{j}}.$
If $A$ and $B$ are both normal, then $\Delta(A,B)=0$, and we obtain $\operatorname{d}(A,B)=\operatorname{sep}_{\mathrm{F}}(A,B)$.
The second one is a perturbation theorem from [19, Theorem 3.5], and Betcke and Kressner gave a new proof in [20].
Theorem 2 (Invariant Subspace Perturbation Theorem [19]).
@@
If $\delta\equiv\operatorname{sep}(A_{11},A_{22})-\left(\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\right)>0$, and $\left\lVert E_{21}\right\rVert\left(\left\lVert A_{12}\right\rVert+\left\lVert E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}$, then there exist a matrix $R\in\mathbb{C}^{(n-l)\times l}$ satisfying $\left\lVert R\right\rVert\leq 2\delta^{-1}\left\lVert E_{21}\right\rVert$, such that $\operatorname{span}\left\{P_{1}\begin{bmatrix}I\\ R\end{bmatrix}\right\}$ is an invariant subspace of $A+E$.
3 Periodic-PEP and Discrete State Transition Matrix
When a multistep method is applied to an LPTV system eq.˜1, it gives rise to a linear difference equation with periodic coefficient matrices, referred to as a periodic difference equation. In this section, we will show that this periodic difference equation corresponds to a periodic polynomial eigenvalue problem (periodic-PEP). Furthermore, the periodic-PEP can be transformed into a periodic eigenvalue problem whose coefficient matrix coincides with the discrete STM over a period, capturing the dynamics of the periodic difference equation.
3.1 Derivation of the Periodic-PEP
As in discussion of one-step method, we consider the period $[0,T]$ is partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, In fact, this grid can be extended to an infinite sequence $\left\{t_{i}\right\}_{i\in\mathbb{Z}}$ satisfying $t_{i+p}=t_{i}+T$. This extension defines a periodic time grid. Throughout this paper, quantities that exhibit periodicity with period $p$ on this periodic discrete grid are indexed with a superscript $(i)$ to indicate their cyclic dependence. Hence, the coefficients $(\alpha_{j,i},\beta_{j,i})$ are denoted by $(\alpha_{j}^{(i)},\beta_{j}^{(i)})$, and the step size $\Delta t^{(i)}=t_{i}-t_{i-1}$, and the maximum step $\Delta t=\max_{i=1}^{p}\Delta t^{(i)}$.
$\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)x_{i-d+j}=0,$ (12)
Since $\left(I_{n}-\beta_{d}^{(i)}G^{(i)}\Delta t^{(i)}\right)$ is invertible for sufficiently small step size $\Delta t^{(i)}$, eq.˜12 can be expressed in the form of a $d$-step difference equation with $p$-periodic coefficients:
$x_{i}-\sum_{j=0}^{d-1}A_{j}^{(i)}x_{i-d+j}=0,\quad i=1,\ldots,p,$ (13)
where the coefficient matrices $A_{j}^{(i)}$ for $j=0,\ldots,d-1$ are defined by
$A_{j}^{(i)}=-\left(I_{n}-\beta_{d}^{(i)}G^{(i)}\Delta t^{(i)}\right)^{-1}\left(\alpha_{j}^{(i)}I_{n}-\beta_{j}^{(i)}G^{(i-d+j)}\Delta t^{(i)}\right).$ (14)
In this section, we only use the cyclicity of $A_{j}^{(i)}$ and do not require any of its specific properties from the multistep method.
Next, we define the periodic-PEP corresponding to eq.˜13.
Definition 4 (Periodic Polynomial Eigenvalue Problem).
Let $\{A_{0}^{(i)},A_{1}^{(i)},\ldots,A_{d-1}^{(i)}\}_{i=1}^{p}$ be $p$ sets of coefficient matrices in $\mathbb{R}^{n\times n}$. The periodic polynomial eigenvalue problem (periodic-PEP) seeks scalars $\theta\in\mathbb{C}$ and nonzero vectors $u^{(i)}\in\mathbb{C}^{n}$, $i=1,\ldots,p$, satisfying the periodicity condition $u^{(i+p)}=u^{(i)}$ such that
$u^{(i)}\theta^{d}-\sum_{j=0}^{d-1}A_{j}^{(i)}\,u^{(i-d+j)}\theta^{j}=0,\qquad i=1,\ldots,p.$ (15)
If $(\theta,\{u^{(i)}\}_{i=1}^{p})$ is an eigenpair, then for each $k=0,\ldots,p-1$, $\left(\theta\zeta_{p}^{k},\,\{u^{(i)}\zeta_{p}^{ik-k}\}_{i=1}^{p}\right)$ is also an eigenpair, where $\zeta_{p}=e^{\jmath 2\pi/p}$ is a primitive $p$th root of unity. Among them, we define the periodic eigenvalue of the periodic-PEP as $\lambda=\theta^{p}$, and the periodic eigenvector as the sequence $\{u^{(i)}\}_{i=1}^{p}$ associated with the root $\theta$ whose argument $\arg(\theta)\in[0,2\pi/p)$ is minimal.
For $d=1$, the periodic-PEP reduces to the standard periodic eigenvalue problem discussed in [11]. For $d>1$, the problem takes a polynomial form, which we now illustrate. As an example, when $d=2$, and $p=3$, definition˜4 can be written as:
$\left(\theta^{2}I_{3n}-\begin{bmatrix}&&A_{1}^{(1)}\\ A_{1}^{(2)}&&\\ &A_{1}^{(3)}&\\ \end{bmatrix}\theta-\begin{bmatrix}&A_{0}^{(1)}&\\ &&A_{0}^{(2)}\\ A_{0}^{(3)}&&\\ \end{bmatrix}\right)\begin{bmatrix}u^{(1)}\\ u^{(2)}\\ u^{(3)}\\ \end{bmatrix}=0.$ (16)
This example shows that stacking the $p$ component equations in definition˜4 naturally yields a PEP whose coefficient matrices possess a block-cyclic structure. To express this structure compactly, we introduce the stacking notation

### v2

@@ 4.4 Convergence Analysis
where the block structure is consistent with the decomposition of $\mathcal{Z}(t_{0})$ into principal and parasitic components.
Moreover, under Assumptions (A1)–(A4), there exist constants $C_{1},C_{2}>0$, independent of the periodic time grid, such that
$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{X}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq C_{1}h^{s},\qquad\left\lVert\begin{bmatrix}E_{\mathcal{Y}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq C_{2}.$
Proof 4.5.
We begin by expressing $F_{p,0}$ in the basis $\mathcal{Z}(t_{0})$.
$\begin{aligned} &\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})\\ =&\mathcal{Z}^{-1}(t_{0})\mathcal{Z}(t_{p})\left[\mathcal{Z}^{-1}(t_{p})L^{(p)}\mathcal{Z}(t_{p-1})\right]\cdots\left[\mathcal{Z}^{-1}(t_{1})L^{(1)}\mathcal{Z}(t_{0})\right]\\ =&\mathcal{Z}^{-1}(t_{0})\mathcal{Z}(t_{p})\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)\\ =&\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)\end{aligned}.$
We then define the global error matrix and apply a telescoping sum expansion:
$\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}&E_{\mathcal{X}\mathcal{Y}}\\ E_{\mathcal{Y}\mathcal{X}}&E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\equiv$ $\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)-I_{dn}$
$=$ $\sum_{i=0}^{p-1}(I_{dn}+\Delta_{p})\cdots(I_{dn}+\Delta_{p-i+1})\Delta_{p-i}.$
We now estimate the two block columns separately. For the first block column, the telescoping representation yields
$\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{X}}\end{bmatrix}$= $\sum_{i=0}^{p-1}(I_{dn}+\Delta_{p})\cdots(I_{dn}+\Delta_{p-i})\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{X}}\\ E_{p-i,\mathcal{Y}\mathcal{X}}\end{bmatrix}$.
By Lemma˜4.2, there exists a constant $C>0$ such that
$\left\lVert\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{p-i}\right)\right\rVert_{2}\leq\prod_{k=p-i}^{p}\left\lVert I_{dn}+\Delta_{k}\right\rVert_{2}\leq\prod_{k=p-i}^{p}(1+Ch^{(k)})\leq e^{CT},$
which is uniformly bounded for all $i=1,\ldots,p$.
Therefore, we obtain the estimate
$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{X}}\end{bmatrix}\right\rVert_{2}\leq$ $\sum_{i=1}^{p}e^{CT}\left\lVert\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{X}}\\ E_{p-i,\mathcal{Y}\mathcal{X}}\end{bmatrix}\right\rVert_{2}\leq e^{CT}\sum_{i=0}^{p-1}(h^{(p-k)})^{s+1}\leq e^{CT}h^{s}$
where we used $\sum_{i=1}^{p}h^{(i)}=T$ and $h=\max_{i}h^{(i)}$.
An analogous argument applies to the second block column,
$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{Y}}\\ E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq$ $\sum_{i=1}^{p}e^{CT}\left\lVert\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{Y}}\\ E_{p-i,\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}$
[…]
By Assumption (A4), the cumulative deviation of the step ratios is uniformly bounded, and therefore the right-hand side is bounded by a constant. This completes the proof.
The global difference estimates established above quantify the deviation between the one-period discrete propagator $F_{p,0}$ and its reference counterpart constructed $\Psi(t_{p},t_{0})$. To link it to estimates of eigenpairs difference, we recall classical results from invariant subspace perturbation theorem [22, 3].
Theorem 4.6 (Invariant Subspace Perturbation Theorem [22]).
@@
If $\delta\equiv\operatorname{sep}(A_{11},A_{22})-\left(\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\right)>0$, and $\left\lVert E_{21}\right\rVert\left(\left\lVert A_{12}\right\rVert+\left\lVert E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}$, then there exist a matrix $R\in\mathbb{C}^{(n-l)\times l}$ satisfying $\left\lVert R\right\rVert\leq 2\delta^{-1}\left\lVert E_{21}\right\rVert$, such that $\operatorname{span}\left\{P\begin{bmatrix}I\\ R\end{bmatrix}\right\}$ is an invariant subspace of $A+E$.
Remark 4.7.
The separation between two square matrices $A_{11}$ and $A_{22}$ is defined as
$\operatorname{sep}_{2}(A_{11},A_{22}):=\inf_{\|X\|_{2}=1}\|A_{11}X-XA_{22}\|_{2}.$
If $A_{11}$ and $A_{22}$ are both normal, then the separation coincides with the spectral gap between their spectra.